# GraspDiffusion: Synthesizing Realistic Whole-body Hand-Object Interaction

### v2

@@ 4 Experiments
Figure 7: Processed HICO-DET [10] dataset sample. We jointly extract the joint depth map for the object and human 3D pose, the human skeleton joint map, and semantic segmentation maps.
@@ 4.1 Dataset Construction
To compensate for the lack of realistic, 3D-annotated human-object interaction datasets, we designed an annotation pipeline through which we leveraged previous interaction datasets [20, 19, 10, 45, 30] to construct a pseudo-3D interaction dataset. Specifically, we utilized the human-object interaction images from HICO-DET [10] and V-COCO [20], which contain a large variety of possible interactions and annotations for the human body and object type.
For both datasets, we first filtered the images so that each image included at least one visible human with a reasonable screen size, along with at least one identifiable hand. For datasets like HICO-DET, multiple humans within an image might be interacting with multiple objects, making it necessary to correctly identify the main object of interaction. We use the BLIP-2 language model [42] to perform a Visual Question Answering task, which outputs an object type from the input image. Using the object text, we use GroundingDINO [46] to get the object location and detect the segmentation map of the object, and employ a state-of-the-art depth estimation model [33] to estimate the depth map of the object. Note that for V-COCO, we use the original annotated information for objects Meanwhile, we also estimate the 3D SMPL-X parameters for the human, using the annotated bounding box and HybrIK [40, 41] 3D human pose and shape estimator. For the human skeleton, we use the annotated Halpe dataset [30] for the HICO-DET dataset and the DWPose estimator [9] for the V-COCO dataset. Among the processed images, we identify images with sufficiently large hands portrayed and reserve them for the hand-object refinement model training. To estimate the 3D MANO parameters, we use the ACR [83] hand pose and shape estimator. To further augment the dataset, we use the BEHAVE interaction dataset [4] that comes with SMPL-X parameters and object 3D models, and processed it in a similar manner to create joint image interaction pairs (image, text, depth map, skeleton, segmentation).
For the hand refinement modules, we process the Dex-YCB dataset [11], the RHD dataset [90] that also comes with 3D annotations and video data. By processing the MANO parameters [63] and the object 3D models, we also acquired image interaction pairs specifically focused on the hand region. To broaden the hand-object interaction distribution, we include a subset of the preprocessed HICO-DET dataset, cropped on the hand-object bounding box, as additional training data for hand refinement. In total, we collected 25K joint interaction pairs of the full human-object interaction scenes to train the scene generation pipeline, and 15K joint interaction pairs of the hand-object interaction scenes to train the hand refinement pipeline.
@@
LDM (finetuned) [62] 41.23 $1.45\times 10^{-2}$ 0.671
ControlNet [84] 32.76 $1.23\times 10^{-2}$ 0.71
Champ [89] 40.63 $2.23\times 10^{-2}$ 0.739
@@
ControlNet [84] 99.38 $7.70\times 10^{-2}$ 58.17
HandRefiner [50] 92.48 $7.11\times 10^{-2}$ 61.45
Affordance Diffusion [82] - - 65.69
@@ 4.2 Implementation Details
For the full-body grasping pipeline, we train the body-pose diffusion model on the GRAB [68] dataset, which captures the full-bodied 3D SMPL-X interaction sequences with various objects. After downsampling the motion sequences of GRAB to 30 fps, we collect all frames that has more than 40 contacting vertices between the object and the subject’s right hand, to collect feasible grasps for the training. We then adopt the Adam optimizer [35] with a learning rate of $5\times 10^{-4}$. We train the model with batch size of 2,048 for 50k steps, using 2 RTX 6000 GPUS. For the diffusion schedule, we adapt a cosine noise schedule with $T=1000$.
To implement the scene generation pipeline, we train the two modules with the aforementioned custom datasets. Employing the Stable Diffusion v1.5 [62] as a base model with parameters frozen, we train the conditional modules with a constant learning rate of $10^{-4}$, for 200 epochs on four A100 GPUS which costs approximately 28 hours. For inference, we used a linear multistep scheduler [31] with 30 inference steps using a classifier-free guidance [26] of 3.5. We also support inference using personalized Stable Diffusion models other than the Stable Diffusion v1.5 model used during training, and display results with different models in Fig. 9.
@@ 4.3 Quantitative Results
Full-body generation quality To assess the generation quality, we adopt Frechet Inception Distance (FID) [25] and Kernel Inception Distance (KID) [5]. We compare our results with three baseline models: (1) a finetuned Stable Diffusion v1.5 model [62] conditioned only by the text description; (2) ControlNet [84] with multiple control input; (3) Champ [89], a human image animation method that uses SMPL-X sequences. For Champ, we separately generate a human image as reference and control the body pose of the reference image using Champ’s guidance encoders, without using its motion module. We randomly select 5K images from the training set from our novel human-object dataset for comparison. In addition, to assess the alignment between the intended text prompt and the generated image’s interaction context, we use CLIPScore [24] as an additional evaluation metric. As shown in Table 1, our method can improve image quality and prompt alignment in generating images with human-object interaction.
@@
GOAL [69] 0.461 0.504 6.135
FLEX [70] 0.540 0.252 7.794
COOP [87] 0.841 0.239 4.679
@@
Hand-grasp generation quality We also assess the quality of hand-centric images from the hand refinement pipeline, based on both image quality and plausible hand-object pose. To measure instances of successful hand-object contact, we adopt the contact evaluation setup in Affordance Diffusion [82] and utilize a widely used hand-object detector [64] to measure the object’s contact status. We compare our results with three baseline models : (1) a depth-based ControlNet, (2) HandRefiner [50], and (3) Affordance Diffusion [82]. We evaluated on a subset of the DexYCB dataset [11]. We report the results in Table 2, which shows that our method is capable of outperforming previous methods in creating hand-object images with accurate contact.
3D pose evaluation To evaluate the plausibility of grasping poses for different objects and positions, we constructed a test set of unseen objects distributed far from the original range of the training dataset. We choose 10 novel 3D objects from Dex-YCB [11] and 10 human body shapes, and for each pair, we position the object at 64 random 3D positions, relative to the human body’s pelvis joint position. Specifically, the x-coordinate (the horizontal position in our paper) ranges from -0.5m to 0.5m, the y-coordinate (the vertical position in our paper) ranges from -0.8m to 0.8m, and the z-coordinate (the direction where the human model is facing) ranges from 0.0m to 0.8m, with the pelvis joint position as its origin.
For pose evaluation, we utilize VPoser [56] to measure the L2 loss of vertex reconstruction from the body poses as a pose-valid error, given that an implausible body pose will result in a higher pose-valid error. We compare our approach with two prior methods that are trained on the GRAB dataset [68] and support generating full-body grasps for different object translations; GOAL [69] and COOP [87]. For GOAL, we only evaluate the grasping pose generation with optimization (GNet) and set the x-coordinate of the object translation to 0 due to the fact that GOAL does not work when the objects are out of distribution in the horizontal plane [70]. We present the results in Table  3. The results demonstrate that our model is capable of generating authentic grasping poses for objects in various positions.
@@
User Study We conducted a user study to measure the perceptual quality and geometric consistency of our pipeline. We asked 28 participants to compare images that were generated based on the same 3D grasping pose and object, one generated using multiple ControlNets [84] and HandRefiner [50] and the other using GraspDiffusion. The participants were asked two types of questions: (a) which image is more realistic and plausible compared to the other and (b) given the original grasping information, which one follows faithfully to the grasping context. In total, 92.4% of the votes preferred our method over the baseline on plausibility, while 96.4% preferred based on following the given context.
@@
Figure 9: Example results from different models. We display generation results from our pipeline with the same object and body pose, but with different personalized Stable Diffusion models that were acquired from CivitAI [1] and Huggingface [2].
@@ 4.4 Applications
By utilizing a 3D full-body grasping model generated from a simple object input, GraspDiffusion can provide a practical solution for AI practitioners who intend to use generative AI for their artwork such as advertisements, illustrations, and comic books. To alleviate the requirement of a 3D object mesh model, our pipeline also supports using 3D reconstruction models [71, 77, 49] that can recover 3D mesh models from a single image input. Moreover, our pipeline can support various personalized image domains, including (but not limited to) realistic, anime, pixel art style, and more. In Figure 9 we present results from the same 3D grasping pose, using diverse personalized text-to-image models to support different art styles and backgrounds. Further examples of such use cases will be provided in the Supplementary Material.
@@ 5 Discussions and Limitations
While GraspDiffusion can produce humans with detailed finger articulation and accurate object interaction, several samples exhibit discrepancies between the body’s skin texture and the refined hand’s skin texture. We account this issue to the shortage of balanced, high-quality data samples during training, and opt to construct additional interaction samples to facilitate high-quality generation. In the future, we aspire to extend our pipeline toward scene-level generation that involves interaction between multiple humans and objects, and to better define interactions in the form of user-controllable text prompts. We also look forward to synthesize zero-shot interaction motions by leveraging image-to-video diffusion models [6, 81, 22, 75].
@@ 6 Conclusion
We present an image generation pipeline that is the first to explicitly target realistic human-object interaction. The resulting images exhibit both explicit (hand-object contact, realistic hand grasp) and implicit human-object interaction (human gaze towards the object, body direction), without requiring any auxiliary conditions other than a 3D object mesh and its relative position. The results demonstrate our method’s effectiveness in creating images with plausible hand poses, while preserving the given object’s identity. In the future, we plan to extend our pipeline towards generating various types of interaction (e.g. human-human interaction, specialized hand-object interaction), while further demonstrating the effectiveness of our pipeline in video generation and synthetic dataset creation for interaction detection.

### v3

@@ 4 Experiments
Figure 7: Processed HICO-DET [Chao2017HICODET] dataset sample. We jointly extract the joint depth map for the object and human 3D pose, the human skeleton joint map, and semantic segmentation maps.
@@ 4.1 Dataset Construction
To compensate for the lack of realistic, 3D-annotated human-object interaction datasets, we designed an annotation pipeline through which we leveraged previous interaction datasets [Gupta2015VCOCO, Gkioxari2017DetectingAR, Chao2017HICODET, Li2019HAKEHA, Jin2020WholeBodyHP] to construct a pseudo-3D interaction dataset. Specifically, we utilized the human-object interaction images from HICO-DET [Chao2017HICODET] and V-COCO [Gupta2015VCOCO], which contain a large variety of possible interactions and annotations for the human body and object type.
For both datasets, we first filtered the images so that each image included at least one visible human with a reasonable screen size, along with at least one identifiable hand. For datasets like HICO-DET, multiple humans within an image might be interacting with multiple objects, making it necessary to correctly identify the main object of interaction. We use the BLIP-2 language model [Li2023BLIP2BL] to perform a Visual Question Answering task, which outputs an object type from the input image. Using the object text, we use GroundingDINO [liu2023grounding] to get the object location and detect the segmentation map of the object, and employ a state-of-the-art depth estimation model [ke2023marigold] to estimate the depth map of the object. Note that for V-COCO, we use the original annotated information for objects Meanwhile, we also estimate the 3D SMPL-X parameters for the human, using the annotated bounding box and HybrIK [li2021hybrik, li2023hybrikx] 3D human pose and shape estimator. For the human skeleton, we use the annotated Halpe dataset [Jin2020WholeBodyHP] for the HICO-DET dataset and the DWPose estimator [Cao2018OpenPoseRM] for the V-COCO dataset. Among the processed images, we identify images with sufficiently large hands portrayed and reserve them for the hand-object refinement model training. To estimate the 3D MANO parameters, we use the ACR [Yu2023ACRAC] hand pose and shape estimator. To further augment the dataset, we use the BEHAVE interaction dataset [Bhatnagar2022BEHAVEDA] that comes with SMPL-X parameters and object 3D models, and processed it in a similar manner to create joint image interaction pairs (image, text, depth map, skeleton, segmentation).
For the hand refinement modules, we process the Dex-YCB dataset [Chao2021DexYCBAB], the RHD dataset [Zimmermann_2017_ICCV] that also comes with 3D annotations and video data. By processing the MANO parameters [MANO:SIGGRAPHASIA:2017] and the object 3D models, we also acquired image interaction pairs specifically focused on the hand region. To broaden the hand-object interaction distribution, we include a subset of the preprocessed HICO-DET dataset, cropped on the hand-object bounding box, as additional training data for hand refinement. In total, we collected 25K joint interaction pairs of the full human-object interaction scenes to train the scene generation pipeline, and 15K joint interaction pairs of the hand-object interaction scenes to train the hand refinement pipeline.
@@
LDM (finetuned) [Rombach2021HighResolutionIS] 41.23 $1.45\times 10^{-2}$ 0.671
ControlNet [Zhang_2023_ICCV] 32.76 $1.23\times 10^{-2}$ 0.71
Champ [zhu2024champ] 40.63 $2.23\times 10^{-2}$ 0.739
@@
ControlNet [Zhang_2023_ICCV] 99.38 $7.70\times 10^{-2}$ 58.17
HandRefiner [lu2023handrefiner] 92.48 $7.11\times 10^{-2}$ 61.45
Affordance Diffusion [Ye2023AffordanceDS] - - 65.69
@@ 4.2 Implementation Details
For the full-body grasping pipeline, we train the body-pose diffusion model on the GRAB [Taheri2020GRABAD] dataset, which captures the full-bodied 3D SMPL-X interaction sequences with various objects. After downsampling the motion sequences of GRAB to 30 fps, we collect all frames that has more than 40 contacting vertices between the object and the subject’s right hand, to collect feasible grasps for the training. We then adopt the Adam optimizer [Kingma2014AdamAM] with a learning rate of $5\times 10^{-4}$. We train the model with batch size of 2,048 for 50k steps, using 2 RTX 6000 GPUS. For the diffusion schedule, we adapt a cosine noise schedule with $T=1000$.
To implement the scene generation pipeline, we train the two modules with the aforementioned custom datasets, using 20K of the interaction pairs. Employing the Stable Diffusion v1.5 [Rombach2021HighResolutionIS] as a base model with parameters frozen, we train the conditional modules with a constant learning rate of $10^{-4}$, for 200 epochs on four A100 GPUS which costs approximately 28 hours. For inference, we used a linear multistep scheduler [Karras2022edm] with 30 inference steps using a classifier-free guidance [ho2021classifierfree] of 3.5. We also support inference using personalized Stable Diffusion models other than the Stable Diffusion v1.5 model used during training, and display results with different models in Fig. 9.
@@ 4.3 Quantitative Results
Full-body generation quality To assess the generation quality, we adopt Frechet Inception Distance (FID) [Heusel2017FID] and Kernel Inception Distance (KID) [Binkowski2018KID]. We compare our results with three baseline models: (1) a finetuned Stable Diffusion v1.5 model [Rombach2021HighResolutionIS] conditioned only by the text description; (2) ControlNet [Zhang_2023_ICCV] with multiple control input; (3) Champ [zhu2024champ], a human image animation method that uses SMPL-X sequences. For Champ, we separately generate a human image as reference and control the body pose of the reference image using Champ’s guidance encoders, without using its motion module. We used the 5K testing set from our novel human-object dataset for comparison. In addition, to assess the alignment between the intended text prompt and the generated image’s interaction context, we use CLIPScore [hessel-etal-2021-clipscore] as an additional evaluation metric. As shown in Table 1, our method can improve image quality and prompt alignment in generating images with human-object interaction.
@@
GOAL [Taheri2021GOALG4] 0.461 0.504 6.135
FLEX [Tendulkar2022FLEXFG] 0.540 0.252 7.794
COOP [Zheng_2023_ICCV] 0.841 0.239 4.679
@@
Hand-grasp generation quality We also assess the quality of hand-centric images from the hand refinement pipeline, based on both image quality and plausible hand-object pose. To measure instances of successful hand-object contact, we adopt the contact evaluation setup in Affordance Diffusion [Ye2023AffordanceDS] and utilize a widely used hand-object detector [rong2021frankmocap] to measure the object’s contact status. We compare our results with three baseline models : (1) a depth-based ControlNet, (2) HandRefiner [lu2023handrefiner], and (3) Affordance Diffusion [Ye2023AffordanceDS]. We evaluated on a subset of the DexYCB dataset [Chao2021DexYCBAB]. We report the results in Table 2, which shows that our method is capable of outperforming previous methods in creating hand-object images with accurate contact.
3D pose evaluation To evaluate the plausibility of grasping poses for different objects and positions, we constructed a test set of unseen objects distributed far from the original range of the training dataset. We choose 10 novel 3D objects from Dex-YCB [Chao2021DexYCBAB] and 10 human body shapes, and for each pair, we position the object at 64 random 3D positions, relative to the human body’s pelvis joint position. Specifically, the x-coordinate (the horizontal position in our paper) ranges from -0.5m to 0.5m, the y-coordinate (the vertical position in our paper) ranges from -0.8m to 0.8m, and the z-coordinate (the direction where the human model is facing) ranges from 0.0m to 0.8m, with the pelvis joint position as its origin.
For pose evaluation, we utilize VPoser [SMPL-X:2019] to measure the L2 loss of vertex reconstruction from the body poses as a pose-valid error, given that an implausible body pose will result in a higher pose-valid error. We compare our approach with two prior methods that are trained on the GRAB dataset [Taheri2020GRABAD] and support generating full-body grasps for different object translations; GOAL [Taheri2021GOALG4] and COOP [Zheng_2023_ICCV]. For GOAL, we only evaluate the grasping pose generation with optimization (GNet) and set the x-coordinate of the object translation to 0 due to the fact that GOAL does not work when the objects are out of distribution in the horizontal plane [Tendulkar2022FLEXFG]. We present the results in Table  3. The results demonstrate that our model is capable of generating authentic grasping poses for objects in various positions.
@@
User Study We conducted a user study to measure the perceptual quality and geometric consistency of our pipeline. We asked 28 participants to compare images that were generated based on the same 3D grasping pose and object, one generated using multiple ControlNets [Zhang_2023_ICCV] and HandRefiner [lu2023handrefiner] and the other using GraspDiffusion. The participants were asked two types of questions: (a) which image is more realistic and plausible, and (b) given the original grasping information, which one follows faithfully to the grasping context. In total, 92.4% of the votes preferred our method over the baseline on plausibility, while 96.4% preferred based on following the given context.
@@
Figure 9: Example results from different models. We display generation results from our pipeline with the same object and body pose, but with different personalized Stable Diffusion models that were acquired from CivitAI [CivitAI] and Huggingface [Huggingface].
@@ 4.4 Applications
By utilizing a 3D full-body grasping model generated from an object input, GraspDiffusion can provide a practical solution for AI practitioners who intend to use generative AI for their artwork such as advertisements, illustrations, and comic books. To alleviate the requirement of an object mesh model, our pipeline also supports using 3D reconstruction models [TripoSR2024, wu2024unique3d, Long2023Wonder3DSI] that can recover 3D mesh models from a single image input. Moreover, our pipeline can support various personalized image domains, including (but not limited to) realistic, anime, pixel art style, and more. In Figure 9 we present results from the same 3D grasping pose, using diverse personalized text-to-image models to support different art styles and backgrounds. Further examples are provided in the Supplementary Material.
@@ 5 Discussions and Limitations
While GraspDiffusion can produce humans with detailed finger articulation and accurate object interaction, several samples exhibit discrepancies between the body’s skin texture and the refined hand’s skin texture. We account this issue to the shortage of balanced, high-quality data samples during training, and opt to construct additional interaction samples to facilitate high-quality generation. In the future, we aspire to extend our pipeline toward scene-level generation that involves interaction between multiple humans and objects, and to better define interactions in the form of user-controllable text prompts. We also look forward to synthesize zero-shot interaction motions by leveraging image-to-video diffusion models [Blattmann2023StableVD, yang2024cogvideox, HaCohen2024LTXVideo, kong2024hunyuanvideo].
@@ 6 Conclusion
We present an image generation pipeline that is the first to explicitly target realistic human-object interaction. The resulting images exhibit both explicit (hand-object contact, realistic hand grasp) and implicit human-object interaction (human gaze, body direction), without requiring any auxiliary conditions other than a 3D object mesh and its relative position. The results demonstrate our method’s effectiveness in creating images with plausible hand poses, while preserving the given object’s identity. In the future, we plan to extend our pipeline towards generating various types of interaction (e.g. human-human interaction, specialized hand-object interaction), while further demonstrating the effectiveness of our pipeline in video generation and synthetic dataset creation for interaction detection.
Supplementary Material for GraspDiffusion: Synthesizing Realistic Whole-body Hand-Object Interaction
S1 Model Architecture
For the first stage of our pipeline, we use a diffusion model to synthesize a body pose grasping the input object. The model is trained to predict plausible body parameters (6 DoF body pose, global orientation), conditioned on the object’s relative location $t_{\text{obj}}\in\mathbb{R}^{3}$ and the target hand $c_{\text{left}},c_{\text{right}}\in\{0,1\}$. These conditions are transformed to a conditional embedding $v_{c}$, which is added to the timestep embedding $e_{t}$ and passed to residual blocks within the model, following [pmlr-v139-nichol21a]. We used 3 ResNet blocks for the model, and adopted a cosine noise schedule in training. Additional grasping results are provided in Figure 13, and details on model parameters are provided in Table 5.
Parameter Diffusion Model (Body Pose)
Input Channels 132
Condition Channels 5
Model Channels 1024
ResBlock Number 3
Diffusion Steps 1000
Noise Scheduler Cosine
Table 5: Model architecture for body pose generation diffusion model.
(d) Location
(e) GOAL [Taheri2021GOALG4]
(f) COOP [Zheng_2023_ICCV]
(g) Ours
Figure 10: Grasp synthesis comparison with previous methods.
In Figure 10, we provide an example comparison between our method and previous grasping pose generation methods  [Taheri2021GOALG4, Wu2021SAGASW, Tendulkar2022FLEXFG, Zheng_2023_ICCV]. When given a object with its location relative to the human body (left row), GOAL [Taheri2021GOALG4] and SAGA [Wu2021SAGASW] tend to create distorted poses when the object is far away from the human, as they assume it to be in the same horizontal xy-plane. FLEX assumes a world-centric coordinate system, which leads to pose ambiguity for our scenario. While COOP has a similar objective, it focuses on various object heights, and requires an extensive test-time optimization of 5 different loss terms. We are the first to utilize a lightweight diffusion model in synthesizing body grasping poses.
(a) Object Image
(b) 3D Mesh
(c) Result image
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 11: Synthesis results from a single image. We first synthesize a 3D Mesh from the image using TripoSR [TripoSR2024], InstantMesh [xu2024instantmesh], Real3D [jiang2024real3d] and subsequently used the mesh as input.
Parameter Conditional Encoder(s)
Input Channels $3\times 64$
Output Channels $[320,640,1280,1280]$
ResBlock Number 2
Kernel Size 1
Feature Weight (Body) $[1.0,0.6,1.0]$
Feature Weight (Hand) $[1.0,0.6,1.0]$
Parameter Attention Injection
Human Strength 0.2
Object Strength 1.8
Negative Object Strength -9.0
Weight coefficient $(w^{{}^{\prime}})$ 0.4
Table 6: Model architecture for scene generation models, and inference parameters for attention injection.
Refer to caption Refer to caption Refer to caption
Figure 12: Failure cases for GraspDiffusion.
For the second stage of our pipeline, we use a diffusion model based on the Latent Diffusion [Rombach2021HighResolutionIS] architecture, and attach encoders [Mou2023T2IAdapterLA] that receives spatial features from the synthesized body pose. Specifically, we first provide three spatial conditions from the full-body grasping pose; the human skeleton projection, joint depth map, and the occluded object with ambient lighting $([s^{i},d^{i},o^{i}])$. Then we further refine the hand-object region by providing similar spatial conditions, but centered on the hand region $([s_{h}^{i},d_{h}^{i},o_{h}^{i}])$. For training, we only train the conditional encoders and fix the parameters for the original Stable Diffusion model, encouraging the encoders to be used with other diffusion models finetuned from Stable Diffusion, accounting to our pipeline’s style flexibility.
Object: an axe
Object: a red mug cup
Object: a red plastic bowl
Object: a black wineglass
Object: a black sword
Object: a coffee cup
Object: a dagger
Object: a tomato soup can
Object: a mustard bottle
Object: a baseball bat
Object: a can of spam
Object: a banana
Figure 13: More results on synthesized images from a given object. Objects were gathered from the DexYCB dataset [Chao2021DexYCBAB] and SketchFab [SketchFab]
During inference, we further control the interaction by rectifying the cross-attention maps for the human and object. For the segmentation masks from the body pose (human : $m^{i}$, object : $m^{i}_{o}$, negative object : $m^{i}_{no}$), we assign different levels of strength for each maps to create an input attention matrix $A\in\mathbf{R}^{N_{i}\times N_{t}}$, where $N_{i}$ and $N_{t}$ are the number of image and text tokens. We assign a higher weight for the object masks due to their regional size differences. We then edit the cross attention layers so that it computes the output as $\text{softmax}(\frac{QK^{T}+wA}{\sqrt{d_{k}}})V$, where $Q,K,V$ are the query, key and value embeddings, $d_{k}$ is the dimensionality of $Q$ and $K$, and $w$ is a scalar weight that controls the total strength of user input attention. This encourages the image tokens in the segmented regions to adhere more to the corresponding text tokens, ensuring that the interaction captured by the body pose is well maintained. Following [balaji2022eDiff-I], we calculate $w$ as
$w=w^{{}^{\prime}}\cdot\text{log}(1+\sigma)\cdot\text{max}(QK^{T})$
where $w^{{}^{\prime}}$ is a user defined scalar. Details on model parameters and inference are provided in Table 6. Note that for the feature weights, we assigned a relatively low rate for the joint depth map, to ensure the result image doesn’t overfits to the SMPLX [SMPL-X:2019] mesh’s outline.
S2 Additional Results
We display failure cases for our pipeline in Figure. 12, where We note some failure cases, where the refined hand stands out from the image (left row), the hand shape tends to be uncanny (middle row), or where the complex object texture is not correctly preserved within the image (right row).
We also provide additional results for realistic, full-bodied human object interaction image generation in Figure. 13. We display that our model is capable of producing images of realistic humans interacting with the given object, with high diversity over human identity, body pose, camera angle, background, and other relevant scene context. The results demonstrate our pipeline’s capability in creating realistic grasps for unseen objects.